OpenStudy (nightangel1459):

The equation of circle having a diameter with endpoints (-3, 1) and (5, 7) is A: (x - 4)² + (y - 3)² = 25 B: (x - 1)² + (y - 4)² = 25 C: (x - 1)² + (y - 4)² = 100

OpenStudy (anonymous):

 find their midpoint, and thus the center of the circle and use that in the circle formula

OpenStudy (anonymous):

\(\textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -3}}\quad ,&{\color{blue}{ 1}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ 7}}) \end{array}\qquad % coordinates of midpoint \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right) \\ \quad \\ \quad \\ (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\)

OpenStudy (anonymous):

hmmm ohh.. you'd need the radius, of course, well, just find it by getting distance formula, that'd give you the diameter, the radius is half that

OpenStudy (nightangel1459):

okay, thanks :)

OpenStudy (anonymous):

\(\textit{distance between 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -3}}\quad ,&{\color{blue}{ 1}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ 7}}) \end{array}\qquad % coordinates of midpoint d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

yw